40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 9 23:17:47 2010, kiwamu, Update, SUS, beam pointing has been done 
    Reply  Fri Dec 10 09:58:57 2010, Aidan, Update, SUS, beam pointing has been done 
       Reply  Sat Dec 11 00:41:55 2010, kiwamu, Update, IOO, alignment of IP_ANG is done 
          Reply  Tue Dec 14 01:04:23 2010, kiwamu, Update, SUS, alignment of ITMs and PRM done 
             Reply  Tue Dec 14 15:15:19 2010, kiwamu, Update, SUS, put ETMY and aligned the beam 
                Reply  Tue Dec 14 22:42:56 2010, kiwamu, Update, SUS, installed the last TT . And some more progresses 
                   Reply  Thu Dec 16 01:23:43 2010, kiwamu, Update, SUS, SRM OSEMs installed DSC_2757_ss.jpg
Message ID: 4055     Entry time: Tue Dec 14 22:42:56 2010     In reply to: 4054     Reply to this: 4063
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: installed the last TT . And some more progresses 

  [Koji, Oasmu and Kiwamu]

We made the following progresses :

   (1) installation of the last TT called SR2.

 (2) fixing an earthquake stop issue on the BS

  (3) fixing a clipping of beams at the dark port


(earthquake stops on the BS)

 - When we were aligning the BS, we found that the BS showed funny behaviors.

For example a kick on LR didn't shake LRSEN, and a big DC angular offset (~ 8 in the medm screen ) was needed to keep a horizontal beam axis in the reflection.

We checked all the earthquake stops and found two suspicious earthquake stops at right bottom side.

They looked like slightly touching the BS. So we moved them further away from the BS.

Then the problem had gone.  We doublechecked the health by kicking, applying an angular offset  and so on.

 

(beam clipping)

We found transmitted/reflected beams from the BS to the dark port was clipped at the BS tower.

We moved the BS tower by ~1cm and realigned the rotation of the BS tower.

We also found the beam spots on the two TTs, PR2 and PR3, were offcentered. Especially the beam spot on PR3 was almost on the edge of the mirror.

Probably this was because we touched PZT1 when aligning the beam to the Y end.

So it means, it is not a good idea to align the beam to the end only by steering PZT1. We should use PR2 and PR3 as well when we align the beam to the Y end.

We realigned them such that the beam hits the center of PR2, PR3 and ETMY.

ELOG V3.1.3-